[figure: 9.7 cm]

SUPERCONDUCTIVITY AND NORMAL-STATE BEHAVIOR OF

CO-DOPED (Y_{1−x}Gd_{x})Ba_{2}(Cu_{0.98}Zn_{0.02})_{3}O_{7−δ}

E. Burzo, I.Gr. Deac, V. Pop, R. Tetean Babes-Bolyai University, Faculty of Physics, RO 3400 Cluj-Napoca, Romania

Polycrystalline samples of composition (Y_{1−x}Gd_{x})Ba_{2}(Cu_{0.98}Zn_{0.02})_{3}O_{7−δ} were investigated, for their superconducting and normal-state properties, by using resistivity, ac and dc susceptibility measurements. X-ray powder diffraction has been used to find the lattice parameters of the samples. The doped Zn acts as scatterer and induces a local spin magnetic moment. The interaction of this induced moment with the Gd^{3+} magnetic moment could change the normal behavior of those compounds and magnetic anomalies were shown in dc magnetic susceptibility measurements. Electrical resistance and ac susceptibility data indicated that T_{c} value decreases as the level of Gd doping increases. The conduction mechanism of electrons as revealed by (T) behaviour seems follow a crossover from purely metallic regime to the localization regime due to weak localization or electron-electron interaction. The electrical and magnetic anomalies above T_{c} are also analyzed taking account the presence of a pseudogap (spin-gap) that opens in the density of states above a characteristic temperature T^{∗} (T^{∗} > T_{c}).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : R.Tetean

Address for correspondence : Babes-Bolyai University, Faculty of Physics, RO-3400 Cluj-Napoca, Romania Tel: ++4064194315 ext:196, Fax: ++4064191906

Email address : rote@phys.ubbcluj.ro